OpenStudy (anonymous):

Given: m∠1=m∠2 and m∠2=75 Prove: m∠1=75 Statement Reason m∠1=m∠2 m∠2=75 m∠1=75

OpenStudy (anonymous):

How about due to the "Friggin OBVIOUS" Theorem

OpenStudy (anonymous):

i get that, i just have to prove it i dont know the terms and such.

OpenStudy (anonymous):

ive always been terrible at proofs haha

OpenStudy (ranga):

A = B B = C Therefore, A = C due to transitivity.

OpenStudy (anonymous):

I can't believe they still make students go through stupid stuff like this. Really makes kids hate math. To be technical, I suppose you could say the reason is the Transitive Property of Equality http://www.mathwords.com/t/transitive_property.htm

OpenStudy (anonymous):

thanks, i have a couple more if you/re willing to help. these things are screwing my grades up because i can do the problems i just cant do these dumb proofs

OpenStudy (anonymous):

@BangkokGarrett i have to list reasons for all three things listed would the first be given?

OpenStudy (anonymous):

Given Given Law of Transitivity (different teachers might have slightly different names.)

OpenStudy (anonymous):

thanks a ton! how about this one? Given: x + 9 = 13 Prove: x = 4 Statement Reason x+9=13 x=4

OpenStudy (anonymous):

this one needs 2 reasons. sorry for taking your time these things just ruin this class or me and i need to do them to pass and theres 7 of them haha

OpenStudy (ranga):

x + 9 = 13 (given) x = 4 (subtractive property of inequality. 9 is being subtracted from both sides of the equation)

OpenStudy (anonymous):

Given: 2x + y = 70; y = 3x Prove: 2x + 3x = 70 Statement Reason 2x+y=70; y = 3x 2x+3x=70

OpenStudy (anonymous):

thanks @ranga ! if you could help me with a couple more that would be awesome

OpenStudy (ranga):

2x+y=70; y = 3x (given) 2x+3x=70 (substitution property) Please see this link and choose the appropriate property: http://www.basic-mathematics.com/properties-of-equality.html

OpenStudy (anonymous):

Given: ∠A≅∠B Prove: ∠B≅∠A Statement Reason ∠A≅∠B ∠B≅∠A @ranga is this the transistive or substitution property?or neither/

OpenStudy (ranga):

Symmetric Property. See the second item in the link I posted.

OpenStudy (anonymous):

will do

OpenStudy (anonymous):

i read through it i just dont know what iid put for one about a line segment Given: XY is a segment Prove: XY≅XY Statement Reason XY is a segment given XY≅XY

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

these are all just kind of confusing theres 3 left and im done with proofs for the semester/

OpenStudy (ranga):

The very first item in the link.

OpenStudy (ranga):

Reflexive Property.

OpenStudy (anonymous):

Given:15y + 7 = 12 – 20y Prove: y = 1/7 Statement Reason 15y+7=12-20y 35y+7=12 35y=5 y= 1/7

OpenStudy (anonymous):

this has 4 reasons

OpenStudy (anonymous):

last one, three reasons thanks a ton you're saving my tail Given: m∠1+m∠2=100 and m∠1=80 Prove: m∠2=20 Statement Reason m∠1+m∠2=100 m∠1=80 80+ m∠2=100 m∠2=20

OpenStudy (ranga):

Statement Reason 15y+7=12-20y Given 35y+7=12 Addition Property 35y=5 Subtraction Property y= 1/7 Division Property

OpenStudy (ranga):

Statement Reason m∠1+m∠2=100 Given m∠1=80 Given 80+ m∠2=100 Substitution Property m∠2=20 Subtraction Property